# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks – the GATTACA Framework

### v2

@@ III-C Reinforcement learning
$q_{\pi}(s,a)=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl{[}R_{a}(s,s^{\prime})+\gamma\sum_{a^{\prime}\in\mathcal{AC}}\pi(s^{\prime},a^{\prime})q_{\pi}(s^{\prime},a^{\prime})\Bigr{]}.$ (1)
Value functions define a partial order over policies and there always exists at least one policy which is better than or equal to all other policies [22]. This policy, denoted $\pi_{*}$, is referred to as the optimal policy. Optimal policies share the same optimal action-value function, denoted $q_{*}$, which is defined as
@@
Optimal policies can be computed with dynamic programming (DP) algorithms, such as policy iteration or value iteration, given a perfect model of the environment in the form of an MDP. DP algorithms are obtained by turning Bellman equations into update rules that improve the approximations of the desired value functions. However, the utility of DP algorithms is limited due to the assumption of a perfect model and its computational expense. Therefore, for large state space environments or cases where the model of the environment is unavailable, Monte Carlo methods or temporal-difference learning (TD) methods are used, see [22].
@@ III-D Q-learning
$Q(S_{t},A_{t})\leftarrow Q(S_{t},A_{t})+\alpha[\,R_{t+1}+\gamma\max_{a\in\mathcal{AC}}Q(S_{t+1},a)-Q(S_{t},A_{t})\!],$
@@
$A_{t}\leftarrow\begin{cases}\operatorname*{arg\,max}_{a\in\mathcal{AC}}Q(S_{t},a)&\textrm{with probability }1-\epsilon\\ \textrm{uniformy random action in }\mathcal{AC}&\textrm{with probability }\epsilon,\\ \end{cases}$
@@ III-E Deep reinforcement learning
$\theta_{t+1}\leftarrow\theta_{t}+\alpha\Big{[}R_{t+1}+\gamma\max_{a\in\mathcal{AC}}Q(S_{t+1},a;\theta_{t})$
$-Q(S_{t},A_{t};\theta_{t})\Big{]}\nabla_{\theta_{t}}Q(S_{t},A_{t};\theta_{t}),$
resembling stochastic gradient descent. It was shown that as the agent explores the environment, the parameterized value function converges to the true $Q$-function, see, e.g. [22].
@@ III-E1 Double Deep Q-Network (DDQN)
Both standard Q-learning and DQN use the same values for both selecting and evaluating actions, which can lead to choosing actions based on overestimated values [23]. To address this issue, DDQN introduces two networks with separate sets of weights, $\theta$ and $\theta^{\prime}$, where the former is used to predict actions and the latter to evaluate the greedy policy. They are trained by assigning each experience randomly to update one of the networks. DDQN uses the following DDQN target:
@@ III-E2 Branching Dueling Q-Network (BDQ)
BDQs are designed to address complex and high-dimensional action spaces; they enhance the scalability and sample efficiency of RL algorithms in complex scenarios. BDQs can be thought of as an adaptation of the dueling network architecture [24] into the action branching architecture. A BDQ uses two separate ANNs, i.e. the target network $Q^{-}$ for evaluation and the controller network $Q$ for selection of actions. Furthermore, instead of using a single output layer for all actions, BDQ introduces multiple branches, each responsible for a specific subset of actions. This decomposition allows for parallel processing and shared representations, making it more efficient for the agent to learn and execute tasks with a diverse action space. BDQs avoid overestimating Q-values, can more rapidly identify action redundancies, and generalise more efficiently by learning a general Q-value that is shared across many similar actions.
@@ IV Control problem formulation
Often biologists are interested not in reaching a single, fully specified state, but rather a set of states in which a specified subset of genes attains a certain configuration of expression levels, which remains constant afterward. For example, they may be interested in achieving any state in which apoptosis occurs. Therefore, we assume that the target is given as a combination of desired values for a subset of genes rather than a fully specified attractor as in [7]. We define an attractor-target control strategy as follows.
@@
Figure 1: Schematic illustration of the concept behind Attractor-Target Control. The source attractor $\mathcal{A}_{S}$ (blue dot) is not aligned with the target configuration (target gene expression profile for two of the four genes is encoded with colours: olivegreen - OFF, aquamarine - ON). The target can be any of the attractors (red dots) containing a state that is aligned with the target configuration. Transitions that follow the original network dynamics are shown with dashed arrows. Red lightnings indicate interventions (control actions) and red arrows indicate the change of state due to interventions. The depicted control strategy is of length 3, which from $\mathcal{A}_{S}$ traverses two intermediate attractor states (green dots) to finally reach one of the target attractors. Ellipses represent basins of attraction of the respective attractors on the control trajectory (path).
@@
The concept of the attractor-target control strategy is schematically illustrated in Figure 1 and an example is provided in Appendix -A. Notice that, according to the above definition, we aim to guide the network dynamics toward any state within a subset of (pseudo-)attractor states that match the target configuration, rather than a specific (pseudo-)attractor. This definition is more general than the definition of the ‘Attractor-Based Control Strategy’ introduced in [7], where the system is expected to reach a specific attractor. The current definition of the ‘Attractor-Target Control Strategy’ aligns better with contemporary biological practice.
We now define the attractor-target control problem, which is a variation of the ‘target-based sequential instantaneous control (ASI) problem defined for BNs in [25]. An exact ‘divide-and-conquer’-type algorithm for solving the ASI problem for BNs is implemented in the software tool CABEAN [5], which we use as a benchmark in the evaluation of our proposed approach.
@@ V GATTACA: Graph-based Attractor-Target Control Algorithm
We propose a Graph-based Attractor-Target Control Algorithm (GATTACA) computational framework – a GNN-based DRL approach for solving the attractor-target control problem. With cellular reprogramming in mind, we focus on cellular phenotypic functional states, which are observable in experimental practice, and thus allow the DRL agent, which we refer to as the Attractor-Target Graph-based Controller, to take actions only in (pseudo-)attractor states (PA states for short) in agreement with our control problem formulation in Section IV. Since the identification of all attractor states is an NP-hard problem in itself, we improve the pseudo-attractor states identification procedure (PASIP), which was previously established in [7]. The improved PASIP and the Attractor-Target Graph-based Controller constitute the two main components of the GATTACA framework. The source code of our framework together with the trained models are available at [26].
@@ V-A Improved PASIP (iPASIP)
We observed that the original PASIP method in [7] can be impeded by the presence of large attractors in a BN model. For example, the T-LGL network of [27] has an attractor with more than 1000 states. Of course, if the existence of a large attractor is foreseen and a rough estimate of its size is available, the PASIP $k_{1}$ and $k_{2}$ hyperparameters could be adjusted appropriately based on Theorem .1 presented in Appendix -C, which was originally proved in [7]. However, this would imply the necessity of longer simulations of the BN model, which are computationally expensive. If the existence of a large attractor is not anticipated and the hyper-parameters are not adjusted, the original PASIP is prone to getting trapped in such an attractor while waiting for one of the attractor states to demonstrate a pronounced frequency of re-visits. To handle such cases effectively, we modify PASIP by adding additional checkpoints that enable the identification of a representative state within a large attractor. The improved PASIP (iPASIP) consists of steps executed in two phases: (I) environment pre-processing and (II) DRL agent training. It extends the original PASIP from [7] by incorporating Steps I-2 and II-3 as additional checkpoints.
@@ V-B Attractor-Target Graph-based Controller (ATGC)
where $v$ is the current node, $N(v)$ is the set of neighbours of $v$, and $h$ is the graph convolutional kernel with learnable parameters $\Theta$. In our case, $h$ consists of a neural network with a 64-neuron deep layer and ReLu activations. Here $h(x||y)$ means conditioning-by-intervention, that is the distribution of $x$ with $y$ set to some fixed value. Its role is to reduce spurious correlations present in the graph; for more details, we refer to [28].
@@
Finally, to allow simultaneous perturbation of a combination of genes within a DRL action as in [7], the MLP is followed by a BDQ network architecture [29] [^1] An architecture consisting of the MLP followed by a DDQN was originally proposed in [30]. In [7], we replaced the DDQN with the BDQ architecture to allow multiple perturbations of genes within a single DRL action. For further details, refer to [7].. Each of the five action branches and the state value layer in Figure 2 is a three-layer MLP with 256, 512, and 512 neurons. The idea of including the state value layer is to create the internal representation of the current state, so that it can later be included in the calculation of the Q function of the state-action pair. We want all of the branches to share this common representation, as the value of each action changes depending on the state. We consider five action branches as this is the maximum number of nodes in our study that can be perturbed simultaneously within a single DRL action.
@@ VI-A Case Studies: Models and Their Target Configurations
TABLE I: BN models of real-world biological networks with their target configurations used for the performance evaluation of the GATTACA framework. The BD Id column provides model identification numbers in the BioDivine repository [31].
@@
The logical model for the prediction of bladder cancer cells to become invasive was originally introduced in [32] with the purpose of understanding how genetic alterations combine to promote cancer tumorigenesis. Specifically, the focus is on the relationships between the alterations, i.e. mutual exclusivity and co-occurrence. Using statistical analysis on tumour datasets and a logical modelling approach, the study explores patterns of genetic alterations that affect growth factor receptor signalling, cell cycle regulation, and apoptosis. The model confirms conditions where genetic alterations drive tumour progression and suggests that additional mutations, such as PIK3CA or p21CIP deletion, may enhance invasiveness.
@@
The Mitogen-Activated Protein Kinase (MAPK) network is a system of tightly interconnected signalling pathways involved in diverse cellular processes, such as cell cycle, differentiation, survival, and apoptosis. In [33], a comprehensive and generic reaction map of the MAPK signalling network is constructed based on an extensive analysis of published data. To investigate MAPK responses to different stimuli and understand their involvement in cell fate decisions in urinary bladder cancer, specifically the balance between cell proliferation and cell death, a logical model comprising 53 nodes is constructed, incorporating the most crucial components and interactions. The Boolean model is verified to be globally consistent with published data and it is capable of recapitulating salient dynamical properties of the complex biological network.
@@
T cell large granular lymphocyte (T-LGL) leukemia features a clonal expansion of antigen-primed, competent, cytotoxic T lymphocytes (CTL). To understand the signalling components that are crucial for the survival of CTL in T-LGL leukemia, a T-LGL survival signalling network is constructed in [27]. The network integrates the signalling pathways involved in normal CTL activation and the known deregulations of survival signalling in leukemic T-LGL. The network is translated into a BN model that describes the signalling involved in maintaining the long-term survival of competent CTL in humans. The BN model is capable of reproducing a clinically relevant complex process. In particular, it predicts that the persistence of IL-15 and PDGF is sufficient to reproduce all known deregulations in leukemic T-LGL. For further details, we refer to [27].
@@
A vital process for drought tolerance in plants is examined in [34]. To reduce water loss during drought stress conditions, the phytohormone abscisic acid (ABA) triggers the closure of stomata, i.e. microscopic pores on leaf surfaces through which water loss and carbon dioxide uptake occur. Based on an extensive review of the literature on ABA signal transduction responses underlying stomatal closure, a network of 84 nodes and 156 edges is constructed. This network is then translated into a BN model which predictions agree with 82 out of 99 considered experiments. The model reveals nodes that could be controlled to influence stomatal closure and provides a substantial number of novel predictions that could be subject to wet-lab validation.
@@
Cell growth and proliferation in breast cancer incorporating common signalling pathways is analysed in [35]. To account for the fact that in different cell lines the same combination of drugs can produce different synergistic effects leading to heterogenous responses, the generic model is tailored to represent five breast cancer cell lines. In particular, by integrating information on cell line specific mutations, gene expression profiles, and observed behaviours from drug-perturbed experiments, a BN model for the MCF-7 cell line is proposed. The model predictions on cell-line specific protein activities and drug-response behaviours are in agreement with experimental data.
@@
A comprehensive 71-node Boolean network model was developed to represent the regulatory and signaling pathways controlling CD4+ T helper (Th) cell differentiation [36]. The model captures key transcription factors, cytokine receptors, and intracellular signaling components involved in Th lineage specification. Simulations revealed the emergence of stable states corresponding to canonical Th1, Th2, Th17, and Treg subsets, as well as hybrid cell types that co-express lineage-specific markers, reflecting the context-dependent plasticity of Th cells. Environmental inputs, such as cytokine concentrations, were shown to modulate the attainable cell fates, emphasizing the model’s ability to integrate external cues. A reduced 68-node version of the original network was subsequently used as a case study in [25], focusing on specific environmental conditions to investigate dynamic transitions between Th phenotypes. This work demonstrated the model’s applicability to studying immune cell adaptation in diverse microenvironments and highlighted potential regulatory checkpoints for therapeutic intervention.
@@
Caveolin-1 (CAV1) is a vital scaffold protein heterogeneously expressed in both healthy and malignant tissue. In [38], the role of overexpressed CAV1 in T-cell leukemia is studied. In particular, a 188-node BN model of a CD4+ immune effector T-cell is constructed. The model incorporates various signalling pathways and corresponding protein-to-protein, protein-phosphorylation, and kinase interactions to mimic cellular dynamics and molecular signalling under healthy and immunocompromised conditions. The model was used to predict and examine in silico the heterogeneous effects and mechanisms of CAV1. Specifically, to verify the hypothesis that the role of CAV1 in immune synapse formation contributes to immune regulation during leukemic progression. It was verified that the local interactions built into the BN model were able to accurately mimic complex phenomena observed in the laboratory.

### v3

@@ III-C Reinforcement learning
$q_{\pi}(s,a)=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl[R_{a}(s,s^{\prime})\\ +\gamma\sum_{a^{\prime}\in\mathcal{AC}}\pi(s^{\prime},a^{\prime})q_{\pi}(s^{\prime},a^{\prime})\Bigr].$ (1)
Value functions define a partial order over policies and there always exists at least one policy which is better than or equal to all other policies [SB18]. This policy, denoted $\pi_{*}$, is referred to as the optimal policy. Optimal policies share the same optimal action-value function, denoted $q_{*}$, which is defined as
@@
Optimal policies can be computed with dynamic programming (DP) algorithms, such as policy iteration or value iteration, given a perfect model of the environment in the form of an MDP. DP algorithms are obtained by turning Bellman equations into update rules that improve the approximations of the desired value functions. However, the utility of DP algorithms is limited due to the assumption of a perfect model and its computational expense. Therefore, for large state space environments or cases where the model of the environment is unavailable, Monte Carlo methods or temporal-difference learning (TD) methods are used, see [SB18].
@@ III-D Q-learning
$Q(S_{t},A_{t})\leftarrow Q(S_{t},A_{t})+\alpha[\,R_{t+1}+\gamma\max_{a\in\mathcal{AC}}Q(S_{t+1},a)$
$-Q(S_{t},A_{t})\!],$
@@
[…]
@@ III-E Deep reinforcement learning
$\theta_{t+1}\leftarrow\theta_{t}+\alpha\Big[R_{t+1}+\gamma\max_{a\in\mathcal{AC}}Q(S_{t+1},a;\theta_{t})\\ -Q(S_{t},A_{t};\theta_{t})\Big]\nabla_{\theta_{t}}Q(S_{t},A_{t};\theta_{t}),$
resembling stochastic gradient descent. It was shown that as the agent explores the environment, the parameterized value function converges to the true $Q$-function, see, e.g. [SB18].
@@ III-E1 Double Deep Q-Network (DDQN)
Both standard Q-learning and DQN use the same values for both selecting and evaluating actions, which can lead to choosing actions based on overestimated values [HGS16]. To address this issue, DDQN introduces two networks with separate sets of weights, $\theta$ and $\theta^{\prime}$, where the former is used to predict actions and the latter to evaluate the greedy policy. They are trained by assigning each experience randomly to update one of the networks. DDQN uses the following DDQN target:
@@ III-E2 Branching Dueling Q-Network (BDQ)
BDQs are designed to address complex and high-dimensional action spaces; they enhance the scalability and sample efficiency of RL algorithms in complex scenarios. BDQs can be thought of as an adaptation of the dueling network architecture [WSHH+16] into the action branching architecture. A BDQ uses two separate ANNs, i.e. the target network $Q^{-}$ for evaluation and the controller network $Q$ for selection of actions. Furthermore, instead of using a single output layer for all actions, BDQ introduces multiple branches, each responsible for a specific subset of actions. This decomposition allows for parallel processing and shared representations, making it more efficient for the agent to learn and execute tasks with a diverse action space. BDQs avoid overestimating Q-values, can more rapidly identify action redundancies, and generalise more efficiently by learning a general Q-value that is shared across many similar actions.
@@ IV Control problem formulation
Often biologists are interested not in reaching a single, fully specified state, but rather a set of states in which a specified subset of genes attains a certain configuration of expression levels, which remains constant afterward. For example, they may be interested in achieving any state in which apoptosis occurs. Therefore, we assume that the target is given as a combination of desired values for a subset of genes rather than a fully specified attractor as in [ours_tcs]. We define an attractor-target control strategy as follows.
@@
Figure 1: Schematic illustration of the concept behind Attractor-Target Control. The source attractor $\mathcal{A}_{S}$ (blue dot) of a 4-gene network is not aligned with the target configuration specifying the target expression of two genes (expression profile encoding: olivegreen - gene OFF, aquamarine - gene ON, white - gene expression profile unspecified, i.e. either OFF or ON). The target can be any of the attractors (red dots) containing a state that is aligned with the target configuration. Transitions that follow the original network dynamics are shown with dashed arrows. Red lightnings indicate interventions (control actions) and red arrows indicate the change of state due to interventions. The depicted control strategy is of length 3, which from $\mathcal{A}_{S}$ traverses two intermediate attractor states (green dots) to finally reach one of the target attractors. Ellipses represent basins of attraction of the respective attractors on the control trajectory (path).
@@
The concept of the attractor-target control strategy is schematically illustrated in Figure 1 and an example is provided in Appendix -A. Notice that, according to the above definition, we aim to guide the network dynamics toward any state within a subset of (pseudo-)attractor states that match the target configuration, rather than a specific (pseudo-)attractor. This definition is more general than the definition of the ‘Attractor-Based Control Strategy’ introduced in [ours_tcs], where the system is expected to reach a specific attractor. The current definition of the ‘Attractor-Target Control Strategy’ aligns better with contemporary biological practice.
We now define the attractor-target control problem, which is a variation of the ‘target-based sequential instantaneous control (ASI) problem defined for BNs in [SP20a]. An exact ‘divide-and-conquer’-type algorithm for solving the ASI problem for BNs is implemented in the software tool CABEAN [SP21], which we use as a benchmark in the evaluation of our proposed approach.
@@ V GATTACA: Graph-based Attractor-Target Control Algorithm
We propose a Graph-based Attractor-Target Control Algorithm (GATTACA) computational framework – a GNN-based DRL approach for solving the attractor-target control problem. With cellular reprogramming in mind, we focus on cellular phenotypic functional states, which are observable in experimental practice, and thus allow the DRL agent, which we refer to as the Attractor-Target Graph-based Controller, to take actions only in (pseudo-)attractor states (PA states for short) in agreement with our control problem formulation in Section IV. Since the identification of all attractor states is an NP-hard problem in itself, we improve the pseudo-attractor states identification procedure (PASIP), which was previously established in [ours_tcs]. The improved PASIP and the Attractor-Target Graph-based Controller constitute the two main components of the GATTACA framework. The source code of our framework together with the trained models are available at [GATTACA_git_am].
@@ V-A Improved PASIP (iPASIP)
We observed that the original PASIP method in [ours_tcs] can be fooled by the presence of large attractors in a BN model. For example, the T-LGL network of [tlgl] has an attractor with more than 1000 states. Of course, if the existence of a large attractor is foreseen and a rough estimate of its size is available, the PASIP $k_{1}$ and $k_{2}$ hyperparameters could be adjusted appropriately based on Theorem .1 presented in Appendix -C, which was originally proved in [ours_tcs]. However, this would imply the necessity of longer simulations of the BN model, which are computationally expensive. If the existence of a large attractor is not anticipated and the hyper-parameters are not adjusted, the original PASIP is prone to getting trapped in such an attractor while waiting for one of the attractor states to demonstrate a pronounced frequency of re-visits. To handle such cases effectively, we modify PASIP by adding additional checkpoints that enable the identification of a representative state within a large attractor. The improved PASIP (iPASIP) consists of steps executed in two phases: (I) environment pre-processing and (II) DRL agent training. It extends the original PASIP from [ours_tcs] by incorporating Steps I-2 and II-3 as additional checkpoints.
@@ V-B Attractor-Target Graph-based Controller (ATGC)
where $v$ is the current node, $N(v)$ is the set of neighbours of $v$, and $h$ is the graph convolutional kernel with learnable parameters $\Theta$. In our case, $h$ consists of a neural network with a 64-neuron deep layer and ReLu activations. Here $h(x||y)$ means conditioning-by-intervention, that is the distribution of $x$ with $y$ set to some fixed value. Its role is to reduce spurious correlations present in the graph; for more details, we refer to [Wang2019].
@@
Finally, to allow simultaneous perturbation of a combination of genes within a DRL action as in [ours_tcs], the MLP is followed by a BDQ network architecture [bdq] [^1] An architecture consisting of the MLP followed by a DDQN was originally proposed in [MCSW22]. In [ours_tcs], we replaced the DDQN with the BDQ architecture to allow multiple perturbations of genes within a single DRL action. For further details, refer to [ours_tcs].. Each of the five action branches and the state value layer in Figure 2 is a three-layer MLP with 256, 512, and 512 neurons. The idea of including the state value layer is to create the internal representation of the current state, so that it can later be included in the calculation of the Q function of the state-action pair. We want all of the branches to share this common representation, as the value of each action changes depending on the state. We consider five action branches as this is the maximum number of nodes in our study that can be perturbed simultaneously within a single DRL action.
@@ VI-A Case Studies: Models and Their Target Configurations
TABLE I: BN models of real-world biological networks with their target configurations used for the performance evaluation of the GATTACA framework. The BD Id column provides model identification numbers in the BioDivine repository [BioDivine].
@@
The logical model for the prediction of bladder cancer cells to become invasive was originally introduced in [bladder] with the purpose of understanding how genetic alterations combine to promote cancer tumorigenesis. Specifically, the focus is on the relationships between the alterations, i.e. mutual exclusivity and co-occurrence. Using statistical analysis on tumour datasets and a logical modelling approach, the study explores patterns of genetic alterations that affect growth factor receptor signalling, cell cycle regulation, and apoptosis. The model confirms conditions where genetic alterations drive tumour progression and suggests that additional mutations, such as PIK3CA or p21CIP deletion, may enhance invasiveness.
@@
The Mitogen-Activated Protein Kinase (MAPK) network is a system of tightly interconnected signalling pathways involved in diverse cellular processes, such as cell cycle, differentiation, survival, and apoptosis. In [mapk], a comprehensive and generic reaction map of the MAPK signalling network is constructed based on an extensive analysis of published data. To investigate MAPK responses to different stimuli and understand their involvement in cell fate decisions in urinary bladder cancer, specifically the balance between cell proliferation and cell death, a logical model comprising 53 nodes is constructed, incorporating the most crucial components and interactions. The Boolean model is verified to be globally consistent with published data and it is capable of recapitulating salient dynamical properties of the complex biological network.
@@
T cell large granular lymphocyte (T-LGL) leukemia features a clonal expansion of antigen-primed, competent, cytotoxic T lymphocytes (CTL). To understand the signalling components that are crucial for the survival of CTL in T-LGL leukemia, a T-LGL survival signalling network is constructed in [tlgl]. The network integrates the signalling pathways involved in normal CTL activation and the known deregulations of survival signalling in leukemic T-LGL. The network is translated into a BN model that describes the signalling involved in maintaining the long-term survival of competent CTL in humans. The BN model is capable of reproducing a clinically relevant complex process. In particular, it predicts that the persistence of IL-15 and PDGF is sufficient to reproduce all known deregulations in leukemic T-LGL. For further details, we refer to [tlgl].
@@
A vital process for drought tolerance in plants is examined in [aba]. To reduce water loss during drought stress conditions, the phytohormone abscisic acid (ABA) triggers the closure of stomata, i.e. microscopic pores on leaf surfaces through which water loss and carbon dioxide uptake occur. Based on an extensive review of the literature on ABA signal transduction responses underlying stomatal closure, a network of 84 nodes and 156 edges is constructed. This network is then translated into a BN model which predictions agree with 82 out of 99 considered experiments. The model reveals nodes that could be controlled to influence stomatal closure and provides a substantial number of novel predictions that could be subject to wet-lab validation.
@@
Cell growth and proliferation in breast cancer incorporating common signalling pathways is analysed in [mcf7]. To account for the fact that in different cell lines the same combination of drugs can produce different synergistic effects leading to heterogenous responses, the generic model is tailored to represent five breast cancer cell lines. In particular, by integrating information on cell line specific mutations, gene expression profiles, and observed behaviours from drug-perturbed experiments, a BN model for the MCF-7 cell line is proposed. The model predictions on cell-line specific protein activities and drug-response behaviours are in agreement with experimental data.
@@
A comprehensive 71-node Boolean network model was developed to represent the regulatory and signaling pathways controlling CD4+ T helper (Th) cell differentiation [tdiff]. The model captures key transcription factors, cytokine receptors, and intracellular signaling components involved in Th lineage specification. Simulations revealed the emergence of stable states corresponding to canonical Th1, Th2, Th17, and Treg subsets, as well as hybrid cell types that co-express lineage-specific markers, reflecting the context-dependent plasticity of Th cells. Environmental inputs, such as cytokine concentrations, were shown to modulate the attainable cell fates, emphasizing the model’s ability to integrate external cues. A reduced 68-node version of the original network was subsequently used as a case study in [SP20a], focusing on specific environmental conditions to investigate dynamic transitions between Th phenotypes. This work demonstrated the model’s applicability to studying immune cell adaptation in diverse microenvironments and highlighted potential regulatory checkpoints for therapeutic intervention.
@@
Caveolin-1 (CAV1) is a vital scaffold protein heterogeneously expressed in both healthy and malignant tissue. In [cd4], the role of overexpressed CAV1 in T-cell leukemia is studied. In particular, a 188-node BN model of a CD4+ immune effector T-cell is constructed. The model incorporates various signalling pathways and corresponding protein-to-protein, protein-phosphorylation, and kinase interactions to mimic cellular dynamics and molecular signalling under healthy and immunocompromised conditions. The model was used to predict and examine in silico the heterogeneous effects and mechanisms of CAV1. Specifically, to verify the hypothesis that the role of CAV1 in immune synapse formation contributes to immune regulation during leukemic progression. It was verified that the local interactions built into the BN model were able to accurately mimic complex phenomena observed in the laboratory.